Ticket Number: HEA-NOC/20060517-1 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20060517 12:34 UTC+1 Problem Start: 20060517 12:12 UTC+1
Ticket Closed: 20060811 14:52 UTC+1 Problem End: 20060517 12:57 UTC+1
Listserv services (smtp, web, dns) are not responding to monitoring checks.
Listserv may be unable to send and receive e-mail.
20060517 12:29 LT Ticket opened. The problem is being investigated.
20060517 13:03 LT listserv.heanet.ie was not responding on the console.
The server was rebooted and services are available again. The cause of the
problem is currently being investigated.
20060517 17:59 LT Following the reboot, the server services were
accessible but performance was poor. During troubleshooting, error
messages on a local switch were identified which coincided with the
problem period. The server has two ethernet interfaces, connected to two
switches in failover mode. The switch port connected to the live server
interface was shutdown and the other switch and interface took over. This
has resolved the performance problems and no further errors have been
observed. We are continuing to investigate the root cause of the problem.
20060519 16:57 LT Investigations are ongoing. The errors that were logged
by the Cisco switch are being explored with Cisco support. Performance on
the server has remained stable.
20060602 16:53 LT Further investigation of this problem requires that the
server be connected to two switches again. This will happen in a few weeks
time as part of the next stage of migration of hardware from Brooklawn
House. In the meantime, the server remains stable.
20060811 14:50 LT This issue has not recurred. Closing ticket but will
re-open should problem arise again.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20060517-1